# Stronger Models are NOT Stronger Teachers for Instruction Tuning

### v2

@@ Abstract
Instruction tuning has been widely adopted to ensure large language models (LLMs) follow user instructions effectively. The resulting instruction-following capabilities of LLMs heavily rely on the instruction datasets used for tuning. Recently, synthetic instruction datasets have emerged as an economically viable solution to provide LLMs diverse and high-quality instructions. However, existing approaches typically assume that larger or stronger models are stronger teachers for instruction tuning, and hence simply adopt these models as response generators to the synthetic instructions. In this paper, we challenge this commonly-adopted assumption. Our extensive experiments across five base models and twenty response generators reveal that larger and stronger models are not necessarily stronger teachers of smaller models. We refer to this phenomenon as the Larger Models’ Paradox. We observe that existing metrics cannot precisely predict the effectiveness of response generators since they ignore the compatibility between teachers and base models being fine-tuned. We thus develop a novel metric, named as Compatibility-Adjusted Reward (CAR) to measure the effectiveness of response generators. Our experiments across five base models demonstrate that CAR outperforms almost all baselines.
Stronger Models are NOT Stronger Teachers for Instruction Tuning
[…]
@@ 1 Introduction
To reduce the reliance on human-curated datasets, synthetic datasets generated by LLMs have surfaced as a viable solution Adler et al. (2024). Recent works, such as (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024; Chen et al., 2024), have shown the strong potential of synthetic datasets in instruction tuning. While current research has primarily focused on using LLMs to create large, diverse, and high-quality instructions (Liu et al., 2024b), the selection of appropriate LLMs for generating corresponding responses remains largely unexplored. The common approach is to employ top-performing models (e.g., those leading on benchmarks (Fourrier et al., 2024; Chiang et al., 2024)) for response generation in instruction tuning. For instance, Llama-3.2-3B-Instruct uses responses generated by Llama-3.1-405B-Instruct (i.e., the largest model in Llama-3.1 family) for instruction tuning Meta (2024b). Additionally, most of the existing open synthetic datasets Teknium (2023); Xu et al. (2023a); Ding et al. (2023); Gallego (2023); Chen et al. (2024) depend on expensive, closed-source models like GPT-4 Achiam et al. (2023) and Gemini (Google, 2024) to produce responses.
@@
To further explore the Larger Models’ Paradox, we investigate statistical metrics to reveal potential factors influencing the effectiveness of response generators given different base models. Here, we pose our second research question:
@@
This question is crucial due to the significant computational costs associated with instruction tuning across multiple datasets generated by diverse response generators. Our investigation reveals that existing metrics in alignment data selection, including quality (Dubey et al., 2024), difficulty (Li et al., 2024c), and response length (Liu et al., 2023), fail to consider the compatibility between the base model being fine-tuned and the response generator, thus results in their inability to explain the Larger Models’ Paradox. To bridge this gap, we formulate the task of finding the most effective response generators as a risk-return problem. We solve this by calculating an Compatibility-Adjusted Reward (CAR), where compatibility serves as the risk factor. This compatibility is quantified by the average loss of responses on the base model being fine-tuned, with higher average loss indicating lower compatibility and thus higher risk. Our comparison of the proposed CAR with existing metrics demonstrates that it outperforms all baselines in predicting the effectiveness of response generators.
@@ 2 Related Work
Instruction tuning data selection involves determining which instruction-response pairs to be included in the training dataset and how to sample them (Albalak et al., 2024). The most widely-used metric for selecting instruction data is quality, which is often assessed using LLM evaluators Chen et al. (2023); Liu et al. (2024a), reward models (Dubey et al., 2024; Xu et al., 2024), gradient similarity search (Xia et al., 2024a), or a combination of these methods Cao et al. (2024). Another key metric is difficulty, where higher difficulty is considered more valuable for learning. For instance, Li et al. (2024c) introduces IFD, which measures the instruction-following difficulty of specific instruction-response pairs. Li et al. (2024b) further refines IFD by utilizing GPT-2 for efficient estimation. Approaches like Deita (Liu et al., 2023) consider both quality and difficulty when selecting datasets. Token length is also adopted as a metric, as discussed in (Xia et al., 2024b; Liu et al., 2023).
@@ 3.1 Preliminaries
Supervised fine-tuning (SFT) is widely adopted to enhance instruction-following capabilities of LLMs. The SFT process updates the parameters $\theta$ of a pre-trained language model to minimize the negative log-likelihood loss over the instruction dataset $\mathcal{D}$. The SFT loss can be formally expressed as:
@@ 3.2 Experimental Setup
To evaluate the instruction-following capabilities of the instruction-tuned models, we use two widely-used instruction-following benchmarks: AlpacaEval 2 (AE2) (Li et al., 2023) and Arena-Hard (AH) (Li et al., 2024d). Specifically, AE2 contains 805 representative instructions from real user interactions. AH contains 500 challenging user queries. AE2 and AH use GPT-4-Turbo (1106) and GPT-4-0314 as the baselines to assess the performance of instruction-tuned models, respectively. Both benchmarks compare responses generated by the model of interest with those generated by baselines, and employ GPT evaluators to automatically annotate which response is preferred.
@@ 3.3 Empirical Evaluation
This section evaluates the instruction-following capabilities of models fine-tuned over datasets whose responses are generated by various response generators. By default, we utilize the Magpie-100K dataset as our primary instruction set. Figure 2 provides a comprehensive overview of the AP across different base models and response generators, and the detailed benchmark scores of AE2 and AH are deferred to Table 6 in Appendix B.1.
@@
Our evaluation reveals a counterintuitive finding: increasing the model size of response generators does not necessarily improve the instruction-following capabilities of base models. This finding is universal, evidenced across multiple model families. For example, Gemma-2-9b-it demonstrates superior performance compared to its larger counterpart, Gemma-2-27b-it, in SFT across almost all base models examined. Similar observations are made in other model pairs: Phi-3-Small outperforms Phi-3-Medium, Llama-3.1-70B-Instruct surpasses Llama-3.1-405B-Instruct, Qwen2-7B-Instruct outperforms Qwen2-72B-Instruct, and Qwen2.5-7B-Instruct exceeds Qwen2.5-32B-Instruct. We refer to this finding as the Larger Models Paradox: larger language models, despite their superior performance, may not always generate better responses for fine-tuning smaller language models compared to responses generated by medium-sized models.
@@ 4.2 Baseline Methods
Instruction-following difficulty is another widely-used metric in alignment data selection (Meta, 2024a; Liu et al., 2023; Li et al., 2024c, b; Xu et al., 2024). To assess the difficulty of responses, we employ the following two metrics:
@@
    Instruction Following Difficulty (IFD) (Li et al., 2024c). IFD is defined as:
@@
    where $\text{PPL}(y_{i})$ is the unconditional perplexity of response $y_{i}$. We follow Li et al. (2024b) and employ GPT-2 and the base model respectively, denoted as IFD-GPT2 and IFD-Self.
@@ 4.4 A Compatibility-Aware Metric to Measure Effectiveness
where $r(\mathcal{D}_{i})$ is the average reward measured by the reward model, representing the potential return, and ${L}(\mathcal{D}_{i},\theta)=-\frac{1}{|\mathcal{D}_{i}|}\sum_{y_{i}\in\mathcal{D}_{i}}\log p_{\theta}(y_{i})$ is the average loss for responses in $\mathcal{D}_{i}$ on the base model parameterized by $\theta$. $\beta$ is a tunable parameter that controls the impact of compatibility on the adjusted reward.
@@ B.1 Detailed Benchmark Scores of Instruction-Tuned LLMs
Table 6 details the benchmark scores of AE2 and AH when tuning 5 base models with different response generators. These results complement the Average Performance shown in Figure 2.
@@
B.2 Visualization of baseline methods in measuring the effectiveness of response generators.
@@
Figure 7: PPL-GPT2 and IFD-GPT2.

### v3

# NAACL 2025 Main Conference Stronger Models are Not Always Stronger Teachers for Instruction Tuning

@@ Abstract
Instruction tuning has been widely adopted to ensure large language models (LLMs) follow user instructions effectively. The resulting instruction-following capabilities of LLMs heavily rely on the instruction datasets used for tuning. Recently, synthetic dataset generation has emerged as an economically viable solution to provide LLMs diverse and high-quality instructions. However, existing approaches typically assume that larger or stronger models are stronger teachers for instruction tuning, and hence adopt these models as response generators to distill knowledge for synthetic instructions. In this paper, we challenge this commonly-adopted assumption. Our extensive experiments across five base models and twenty response generators reveal that larger and stronger models are not necessarily stronger teachers of smaller models. We refer to this phenomenon as the Larger Models’ Paradox. We observe that existing metrics cannot precisely predict the effectiveness of response generators since they ignore the compatibility between teachers and base models being fine-tuned. We thus develop a novel metric, named as Compatibility-Adjusted Reward (CAR) to measure the effectiveness of response generators. Our experiments across five base models demonstrate that CAR outperforms almost all baselines.
NAACL 2025 Main Conference
Stronger Models are Not Always Stronger Teachers for Instruction Tuning
Zhangchen Xu    Fengqing Jiang     Luyao Niu    Bill Yuchen Lin     Radha Poovendran University of Washington {zxu9,fqjiang,luyaoniu,byuchen,rp3}@uw.edu
@@ 1 Introduction
To reduce the reliance on human-curated datasets, synthetic datasets generated by LLMs have surfaced as a viable solution Adler et al. (2024). Recent works, such as (Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; Xu et al., 2024; Chen et al., 2024), have shown the strong potential of synthetic datasets in instruction tuning. While current research has primarily focused on using LLMs to create large, diverse, and high-quality instructions (Liu et al., 2024b), the selection of appropriate LLMs for generating corresponding responses remains largely unexplored. The common approach relies on distilling from state-of-the-art models that excel in benchmark evaluations (Fourrier et al., 2024; Chiang et al., 2024) to generate responses for instruction tuning. For instance, Llama-3.2-3B-Instruct uses responses generated by Llama-3.1-405B-Instruct (i.e., the largest model in Llama-3.1 family) for instruction tuning Meta (2024b). Additionally, most of the existing open synthetic datasets Teknium (2023); Xu et al. (2023a); Ding et al. (2023); Gallego (2023); Chen et al. (2024) depend on expensive, closed-source models like GPT-4 Achiam et al. (2023) and Gemini (Google, 2024) to produce responses.
@@
To further explore the Larger Models’ Paradox, we investigate statistical metrics to reveal potential factors influencing the effectiveness of response generators. Here, we pose our second research question:
@@
This question is crucial due to the significant computational costs associated with instruction tuning across multiple datasets generated by diverse response generators. Our investigation reveals that existing metrics in alignment data selection, including quality (Dubey et al., 2024), difficulty (Li et al., 2024d), and response length (Liu et al., 2023), fail to consider the compatibility between the base model being fine-tuned and the response generator, thus results in their inability to explain the Larger Models’ Paradox. To bridge this gap, we formulate the task of finding the most effective response generators as a risk-return problem. We solve this by calculating an Compatibility-Adjusted Reward (CAR), where compatibility serves as the risk factor. This compatibility is quantified by the average loss of responses on the base model being fine-tuned, with higher average loss indicating lower compatibility and thus higher risk. Our comparison of the proposed CAR with existing metrics demonstrates that it outperforms all baselines in predicting the effectiveness of response generators.
@@ 2 Related Work
Instruction tuning data selection involves determining which instruction-response pairs to be included in the training dataset and how to sample them (Albalak et al., 2024). The most widely-used metric for selecting instruction data is quality, which is often assessed using LLM evaluators Chen et al. (2023); Liu et al. (2024a), reward models (Dubey et al., 2024; Xu et al., 2024), gradient similarity search (Xia et al., 2024a), or a combination of these methods Cao et al. (2024). Another key metric is difficulty, where higher difficulty is considered more valuable for learning. For instance, Li et al. (2024d) introduces IFD, which measures the instruction-following difficulty of specific instruction-response pairs. Li et al. (2024c) further refines IFD by utilizing GPT-2 for efficient estimation. Approaches like Deita (Liu et al., 2023) consider both quality and difficulty when selecting datasets. Token length is also adopted as a metric, as discussed in (Xia et al., 2024b; Liu et al., 2023). Selective Reflection-Tuning Li et al. (2024b) approach selects and refines existing instruction-following datasets to address the inconsistency between teacher and student models.
@@ 3.1 Preliminaries
Supervised fine-tuning (SFT) is widely adopted to enhance instruction-following capabilities of LLMs. The SFT updates the parameters $\theta$ of a pre-trained language model to minimize the negative log-likelihood loss over the instruction dataset $\mathcal{D}$. The SFT loss can be formally expressed as:
@@ 3.2 Experimental Setup
To evaluate the instruction-following capabilities of the instruction-tuned models, we use two widely-used instruction-following benchmarks: AlpacaEval 2 (AE2) (Li et al., 2023) and Arena-Hard (AH) (Li et al., 2024e). Specifically, AE2 contains 805 representative instructions from real user interactions. AH contains 500 challenging user queries. AE2 and AH use GPT-4-Turbo (1106) and GPT-4-0314 as the baselines to assess the performance of instruction-tuned models, respectively. Both benchmarks compare responses generated by the model of interest with those generated by baselines, and employ GPT evaluators to automatically annotate which response is preferred.
@@ 3.3 Empirical Evaluation
This section evaluates the instruction-following capabilities of models fine-tuned over datasets whose responses are generated by various response generators. By default, we utilize the Magpie-100K dataset as our primary instruction set. Figure 2 provides a comprehensive overview of the AP across different base models and response generators, and the detailed benchmark scores of AE2 and AH are deferred to Table 7 in Appendix B.1. Evaluations on larger base model (Llama-3.1-8B) with different response generators are presented in Table 6 in Appendix B.2. We analyze the effect of data randomness on average performance in Table 8.
@@
Our evaluation reveals a counterintuitive finding: increasing the model size of response generators does not necessarily improve the instruction-following capabilities of base models within the same model family. This finding is universal, evidenced across multiple model families. For example, Gemma-2-9b-it demonstrates superior performance compared to its larger counterpart, Gemma-2-27b-it, in SFT across almost all base models examined. Similar observations are made in other model pairs: Phi-3-Small outperforms Phi-3-Medium, Llama-3.1-70B-Instruct surpasses Llama-3.1-405B-Instruct, Qwen2-7B-Instruct outperforms Qwen2-72B-Instruct, and Qwen2.5-7B-Instruct exceeds Qwen2.5-32B-Instruct. We refer to this finding as the Larger Models Paradox: larger language models, despite their superior performance, may not always generate better responses for fine-tuning smaller language models within the same model family compared to responses generated by medium-sized models.
@@ 4.2 Baseline Methods
Instruction-following difficulty is another widely-used metric in alignment data selection (Meta, 2024a; Liu et al., 2023; Li et al., 2024d, c; Xu et al., 2024). To assess the difficulty of responses, we employ the following two metrics:
@@
    Instruction Following Difficulty (IFD) (Li et al., 2024d). IFD is defined as:
@@
    where $\text{PPL}(y_{i})$ is the unconditional perplexity of response $y_{i}$. We follow Li et al. (2024c) and employ GPT-2 and the base model respectively, denoted as IFD-GPT2 and IFD-Self.
@@ 4.4 A Compatibility-Aware Metric to Measure Effectiveness
where $r(\mathcal{D}_{i})$ is the average reward measured by the reward model, representing the potential return, and ${L}(\mathcal{D}_{i},\theta)=-\frac{1}{|\mathcal{D}_{i}|}\sum_{y_{i}\in\mathcal{D}_{i}}\log p_{\theta}(y_{i})$ is the average loss for responses in $\mathcal{D}_{i}$ on the base model parameterized by $\theta$. $\beta$ is a tunable parameter that controls the impact of compatibility on the adjusted reward. CAR penalizes the average reward from the reward model with the compatibility risk measured by the loss. This balanced approach enables quantitative assessment of the trade-off between the response quality and compatibility.
@@ B.1 Detailed Benchmark Scores of Instruction-Tuned LLMs
Table 7 details the benchmark scores of AE2 and AH when tuning base models with different response generators. These results complement the Average Performance shown in Figure 2.
B.2 Larger Models’ Paradox in Larger Base Models
We summarize the benchmark scores of AE2 and AH when tuning large base model (Llama-3.1-8B) with diverse response generators in Table 6. We observe that the Larger Models’ Paradox persists when employing the Qwen2.5 and Llama-3.1 model families as response generators. We further demonstrate that the Larger Model’s Paradox is not an effect of data randomness in Table 8.
Table 6: This table presents benchmark scores of AE2 and AH when tuning large base model (Llama-3.1-8B) with diverse response generators. The Larger Models’ Paradox persists when employing the Qwen2.5 and Llama-3.1 model families as response generators.
Base Model Response Generator AE2 LC AE2 WR AH AP
Llama-3.1-8B Qwen2.5-3B-Instruct 11.48 13.85 15.90 13.74
Qwen2.5-7B-Instruct 18.70 20.22 25.90 21.61
Qwen2.5-14B-Instruct 17.50 17.19 28.60 21.10
Qwen2.5-32B-Instruct 16.20 16.42 27.80 20.14
Qwen2.5-72B-Instruct 29.73 32.35 30.90 30.99
Llama-3.1-8B-Instruct 12.62 14.34 15.80 14.25
Llama-3.1-70B-Instruct 14.98 17.74 21.00 17.91
Llama-3.1-405B-Instruct 15.40 17.00 16.50 16.30
Gemma-2-2b-it 17.11 19.64 15.60 17.45
Gemma-2-9b-it 25.74 22.88 23.40 24.00
Gemma-2-27b-it 25.09 24.60 25.40 25.00
B.3 Impact of Data Randomness on Evaluation
We sample 80K instructions from Magpie-100K using different seeds and fine-tuned Llama-3.1-Minitron-4B. The performance of fine-tuned models is shown in Table 8. We observe that the average performance varies by only 2.89%, demonstrating that our evaluation is robust across different instruction samples. This finding underscores the consistency of our evaluation.
@@ B.3 Impact of Data Randomness on Evaluation
B.4 Visualization of baseline methods in measuring the effectiveness of response generators
@@
Figure 7: PPL-GPT2 and IFD-GPT2 of synthetic datasets generated using different response generators.
Table 8: We sample 80K instructions from Magpie-100K using different seeds and fine-tuned Llama-3.1-Minitron-4B with the sampled data. We observe that the average performance varies by only 2.89%, demonstrating that our evaluation is robust across different instruction samples. This finding underscores the consistency of our evaluation.
Instruction Sample AE2 LC AE2 WR AH Average Performance
Magpie-80K (Seed = 42) 14.26 13.54 12.50 13.433
Magpie-80K (Seed = 123) 13.40 12.92 12.80 13.040
Magpie-80K (Seed = 456) 14.77 12.98 11.10 12.950
Magpie-80K (Seed = 789) 13.57 12.79 11.20 12.520
Average 14.00 13.058 11.90 12.986
Standard Deviation 0.634 0.331 0.876 0.375
B.5 Impact of Reward Models on the performance of CAR
We perform ablation analysis on the choice of reward models with a weaker reward model, Skywork-Reward-Llama-3.1-8B, and calculate CAR. The Spearman’s correlations are presented in Table 9. We observe that CAR using the weaker Skywork 8B reward model performs worse compared to using the stronger Skywork 27B reward model, indicating the reliance of CAR on a good performing reward model. However, even with a weaker reward model, CAR outperforms compared with using the reward model alone.
Table 9: Spearman’s correlations when CAR uses different reward models. CAR relies on a good reward model. However, even with a weaker reward model, CAR outperforms compared with using the reward model alone.
Base Model Skywork 8B CAR (Skywork 8B) Skywork 27B CAR (Skywork 27B)
Qwen2-1.5B 0.7895 0.7474 0.8754 0.8842
Gemma 2-2B 0.7982 0.8018 0.8842 0.9000
Qwen2.5-3B 0.7351 0.7386 0.7456 0.8105
Llama-3.1-Minitron-4B 0.8860 0.8912 0.9386 0.9439
Llama-3.2-3B 0.8105 0.8105 0.9088 0.9053